# Uncovering faint lensed gravitational-wave signals and reprioritizing their follow-up analysis using galaxy lensing forecasts with detected counterparts

### v2

@@ 5.2 Re-Ranking the Triggers
To do the re-ranking, for every detected super-threshold lensed multiplet, we make the KDE and compute $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ for each sub-threshold trigger in the list. As the final ranking statistic, we use $\mathcal{B}_{\mathcal{M}_{c}}\times\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ [^7] To some extent, one can see this statistic as the probability ratio of a trigger being lensed compared to unlensed, according to the proximity in chirp mass between the trigger and the super-threshold events and according to the observed time delay..
@@
Figure 4: The posterior overlap statistics ($\mathcal{B}_{\mathcal{M}_{c}}$) (top) and $\mathcal{B}_{\mathcal{M}_{c}}\times\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ (bottom) for the lensed sub-threshold image (orange diamond) and uncorrelated sub-threshold events (black dots). These events are the same scenarios as the one shown in Figs. 3e and 3f, corresponding to the left and right columns, respectively. We see that the presence of a null probability for some time delays discards many triggers. On the left, the genuine counterpart goes from being ranked 83rd out of 123 to being ranked 1st out of 55, while on the right-hand side, we go from rank 65th to 5th, and the list goes from 113 to 62 triggers.
@@
We then carry the procedure out for our 50 lensed multiplets. First, in Fig. 5, we represent the $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ distribution for the lensed and unlensed triggers after having performed a KDE on our results. We see there is some overlap between the two sets of data but the statistic is generally higher for the lensed counterparts than for the unlensed events. Therefore, the lensed counterparts are generally expected to have an improvement in ranking when using this statistic. We then multiply the original ranking statistic $\mathcal{B}_{\mathcal{M}_{c}}$ by $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ and use this as the final ranking for all events. The improvement in ranks and reduction in number of triggers in the lists requiring follow-up is shown in Fig. 6. Some doubles do not have any reduction in the list of triggers because (a) when we see two super-threshold images with different types, there is no region with a zero probability for the time delay, and (b) even when such a region exists for doublets with the same image type, it is very short, generally spanning a few days, compared to the case when we see three super-threshold images, in which the probability of a whole region before or after a certain image is zero. In our examples, we have two scenarios in which the list of triggers is reduced. One is the case where the images have the same image type. The other is when the images have different image types but with a short time delay between them. This happens for low-mass lenses, which means that very long time delays are implausible and some of the triggers with the largest time delays have a negligible probability under the lensed hypothesis due to numerical errors. Therefore, as expected, the method works better when more images are used to constrain the system. Nevertheless, the rank improvement is still effective in all cases considered.
@@
Breaking down the cases of two and three detected lensed images, we can see different effects. When two lensed super-threshold GWs are observed, the trigger list is not significantly improved — the median change is null but its average is a reduction of two triggers — but the rank is. For the latter, on average, the triggers start with rank 66 and end up with rank 5, meaning they go from being in the top 48.5% to being in the top 3.7%. When we observe three super-threshold events, both effects become important, the rank of the genuine counterpart goes from 82 to 9 on average, while the list is reduced by a factor of 2.6, going from 133 to 52 triggers requiring follow-up. So, in this case, on average, we go from counterparts ranked in the top 61.6% initially to them being ranked in the top 17.3% after applying $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$. The difference in effect comes from the difference in characteristics we can build for the different cases, with milder probability differences but regions with zero support for the triples, compared to something more variable for two images but without exclusion regions. Putting all cases together, on average, our method enables one to go from an event ranked 75 out of 133, to an event with rank 7 out of 122, meaning it is generally in the top 10 of a slightly reduced list. So, our method consistently improves the ranking of the genuine counterpart but also directly discards some of the events without requiring more expensive follow-up methods.
@@
It is also worth noting that our approach is computationally cheap compared to running parameter estimation methods on the triggers. The heaviest part is to obtain the initial lensed catalogue from LeR, which takes $\mathcal{O}(hr)$ [^8] Using 8 cores on an Apple M2 Pro CPU.. This operation needs to be done a single time per observation condition as the KDEs required to calculate $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ can all be derived from it. Once we have a catalogue, making the time delay probability distributions and computing the ranking statistics for all the triggers only takes a few minutes.
@@
Finally, let us also stress that, although our approach generally improves the rank of the genuine lensed counterpart and reduces the list of candidates, it does not guarantee that the event we are looking for is top-ranked. So, the safest approach if one does not want to miss the counterpart is still to go through the entire —or at least a significant fraction– of the re-ranked list with more advanced analysis methods. When one wants to do this fully, the list reduction offered by our approach also turns out to be important. Additionally, as mentioned above, our method can also be combined with other methods used to classify or veto triggers, such as imposing a sky overlap between the detected image(s) and the trigger (Wong et al., 2021) or imposing proximity in the chirp masses estimated using by the low-latency searches (Goyal et al., 2023). Another possibility to make lensing searches more efficient is to include the lensing statistics information in the sub-threshold search likelihood itself (Li et al., in prep). We leave the study of combining our approach with others, as well as the application of our method on realistic sub-threshold search results, for future work.
@@ 6 Conclusion and Future Prospects
In this paper, we investigate the use of existing super-threshold lensed gravitational-wave detections to predict the time delays of sub-threshold lensed images of additional images, assuming strong galaxy lensing. Taking the ratio of this probability with that of the time delay under the unlensed hypothesis, one can make a ranking statistic for sub-threshold triggers translating the likelihood of them being a counterpart to the observed super-threshold lensed event. This re-ranking promotes the genuine counterpart as it should be in the higher density region of the lensed distribution. In some cases, we also expect it to reduce the list of triggers requiring follow-up since some time delays have a zero probability under the lensed hypothesis. This effect is due to the fixed ordering of the images as well as our capacity to measure the difference in Morse factors unambiguously in most cases.
@@
To test the effectiveness of the method, we considered a simplified setup consisting of 50 lensed systems made of two or three super-threshold images and, at least, one sub-threshold image. For the super-threshold images, we ran joint parameter estimation to obtain their combined posteriors. To emulate an initial list of candidates, we generate sub-threshold events from a population model, construct posterior distributions following a Gaussian centred on the true values with widths scaled according to the events’ SNR, and compute the posterior overlap statistics between the chirp mass posterior for the unlensed counterpart and the super-threshold images. While not the same as the result from a genuine sub-threshold search system, this somewhat resembles other re-ranking methods proposed in the literature to follow-up on sub-threshold candidates. For every system considered, we compute our ranking statistics based on the measured time delay for the sub-threshold triggers. The final ranking is then the product between the posterior overlap and our statistics. When doing this, we effectively reduce the number of triggers to follow-up, with an average factor improvement of 1.2 but the trigger list is more than halved when three super-threshold images are observed. Moreover, we have improved rankings in all cases. This demonstrates our method is capable of improving the trigger list one would pass to more expensive follow-up searches.
@@
We note that, in principle, if more than one super-threshold image is identified, one could use the joint posteriors for the two first images to make the template bank and evaluate the sky overlap if desired. While this would probably not impact much the template bank itself, the sky overlap condition would probably become more stringent since joint parameter estimation effectively reduces the observed sky area (Janquart et al., 2021a). Moreover, sky overlap is complementary to our approach for reducing strong lensing detection candidates, and therefore using both the ranking based on sky map overlap and on the time delay compatibility would likely further reduce the list of candidates requiring follow-up. Finally, let us mention that if a pair of super-threshold events is detected, predicting and observing a third and/or fourth sub-threshold counterpart would strengthen the case for lensing.
@@
Of course, more testing and development are needed to use the method in realistic scenarios. For example, one could improve the selection of the measured relative lensing parameters to evaluate the $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ ratio, accounting for their full distribution and not their median values. In addition, one could verify the compatibility with other ranking and vetoing methods proposed in the literature, such as the ones used in Wong et al. (2021) and Goyal et al. (2023), for example. Ultimately, these methodologies should be tested on realistic sub-threshold search results, which is beyond the scope of this paper.
@@
In the end, this work shows that one can use the observed relative lensing parameters for a super-threshold lensed doublet or triplet to build a probability distribution for the time delay of a possible sub-threshold counterpart —predicting where such events should be situated in the data stream. Using a simplified setup, we showed that using this distribution to rank sub-threshold triggers is an effective approach because it generally promotes the genuine counterpart and reduces the list of triggers requiring more extensive follow-up. Finally, this method is also complementary to other ranking methods proposed in the literature, such as using the sky overlap between events for example. We hope that our work gives further motivation to the use of the lensing time delay probability distributions in strong lensing searches as a mean to reduce the number of triggers requiring more expensive follow-up analyses.
@@ Acknowledgements
The data underlying this article will be shared in reasonable request to the corresponding authors.
@@ Appendix A Breakdown of the Probability Distributions for Different Sub-cases
In this appendix, we show a case-by-case breakdown of the probability distributions for the different sub-cases explained in Sec. 4 and Table 1.
(a) Total conditioned KDE
(b) Conditioned KDEs of individual sub-cases
@@
(a) Total conditioned KDE
(b) Conditioned KDEs of individual sub-cases
@@
(a) Total conditioned KDE
(b) Conditioned KDEs of individual sub-cases
@@
(a) Total conditioned KDE
(b) Conditioned KDEs of individual sub-cases
@@
(a) Total conditioned KDE
(b) Conditioned KDEs of individual sub-cases
@@
(a) Total conditioned KDE
(b) Conditioned KDEs of individual sub-cases

### v3

@@ 5.2 Re-Ranking the Triggers
To do the re-ranking, for every detected super-threshold lensed multiplet, we make the KDE and compute $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ for each sub-threshold trigger in the list. As the final ranking statistic, we use $\mathcal{B}_{\mathcal{M}_{c}}\times\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$.
@@
Figure 4: The posterior overlap statistics ($\mathcal{B}_{\mathcal{M}_{c}}$) (top) and $\mathcal{B}_{\mathcal{M}_{c}}\times\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ (bottom) for the lensed sub-threshold image (orange diamond) and unrelated sub-threshold events (black dots). These events are the same scenarios as the one shown in Figs. 3e and 3f, corresponding to the left and right columns, respectively. We see that the presence of a null probability for some time delays discards many triggers. On the left, the genuine counterpart goes from being ranked 83rd out of 123 to being ranked 1st out of 55, while on the right-hand side, we go from rank 65th to 5th, and the list goes from 113 to 62 triggers. Note that triggers may not be promoted to the first place because of the statistical nature of our approach.
@@
We then carry the procedure out for our 50 lensed multiplets. First, in Fig. 5, we represent the $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ distribution for the lensed and unlensed triggers after performing a KDE on our results. We see there is some overlap between the two sets of data but the statistic is generally higher for the lensed counterparts than for the unlensed events. Therefore, the lensed counterparts are generally expected to have an improvement in ranking when using this statistic. We then multiply the original ranking statistic $\mathcal{B}_{\mathcal{M}_{c}}$ by $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ and use this as the final ranking for all events. The improvement in ranks and reduction in number of triggers in the lists requiring follow-up is shown in Fig. 6. Some doubles do not have any reduction in the list of triggers because (a) when we see two super-threshold images with different types, there is no region with a zero probability for the time delay, and (b) even when such a region exists for doublets with the same image type, it is very short, generally spanning a few days, compared to the case when we see three super-threshold images, in which the probability of a whole region before or after a certain image is zero. In our examples, we have two scenarios in which the list of triggers is reduced. One is the case where the images have the same image type. The other is when the images have different image types but with a short time delay between them. This happens for low-mass lenses, which means that very long time delays are implausible and some of the triggers with the largest time delays have a negligible probability under the lensed hypothesis due to numerical errors. Therefore, as expected, the method works better when more images are used to constrain the system. Nevertheless, the rank improvement is still effective in all cases considered.
@@
Breaking down the cases of two and three detected lensed images, we can see different effects. When two lensed super-threshold GWs are observed, the trigger list is not significantly reduced: the average reduction is 2, but the ranking of the counterpart is significantly improved. For the latter, on average, the triggers start with rank 66 and end up with rank 5, meaning they go from being in the top 48.5% to being in the top 3.7%. When we observe three super-threshold events, both effects become important, the rank of the genuine counterpart goes from 82 to 9 on average, while the list is reduced by a factor of 2.6, going from 133 to 52 triggers requiring follow-up. So, in this case, on average, we go from counterparts ranked in the top 61.6% initially to them being ranked in the top 17.3% after applying $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$. The difference in effect comes from the difference in characteristics we can build for the different cases, with milder probability differences but regions with zero support for the triples, compared to something more variable for two images but without exclusion regions. Putting all cases together, on average, our method enables one to go from an event ranked 75 out of 133, to an event with rank 7 out of 122, meaning it is generally in the top 10 of a slightly reduced list. So, our method consistently improves the ranking of the genuine counterpart but also directly discards some of the events without requiring more expensive follow-up methods.
@@
It is also worth noting that our approach is computationally cheap compared to running parameter estimation methods on the triggers. The heaviest part is to obtain the initial lensed catalogue from LeR, which takes $\mathcal{O}(hr)$ [^8] Using 8 cores on an Apple M2 Pro CPU.. This operation needs to be done a single time per observation condition as the KDEs required to calculate $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ can all be derived from it. Once we have a sufficiently large catalogue, making the time delay probability distributions and computing the ranking statistics for all the triggers only takes a few minutes.
@@
Finally, let us also stress that, although our approach generally improves the rank of the genuine lensed counterpart and reduces the list of candidates, it does not guarantee that the event we are looking for is top-ranked. So, the safest approach if one does not want to miss the counterpart is still to go through the entire—or at least a significant fraction—of the re-ranked list with more advanced analysis methods. When one wants to do this fully, the list reduction offered by our approach also turns out to be important. Additionally, as mentioned above, our method can also be combined with other methods used to classify or veto triggers, such as imposing a sky overlap between the detected image(s) and the trigger (Wong et al., 2021) or imposing proximity in the chirp masses estimated using by the low-latency searches (Goyal et al., 2023). Another possibility to make lensing searches more efficient is to include the lensing statistics information in the sub-threshold search likelihood itself (Li et al., in prep). We leave the study of combining our approach with others, as well as the application of our method on realistic sub-threshold search results, for future work.
@@ 6 Conclusion and Future Prospects
In this paper, we investigate the use of existing super-threshold lensed gravitational-wave detections to predict the time delays of additional sub-threshold lensed images, assuming strong galaxy lensing. Taking the ratio of this probability with that of the time delay under the unlensed hypothesis, one can make a ranking statistic for sub-threshold triggers translating the likelihood of them being a counterpart to the observed super-threshold lensed event. This re-ranking promotes the genuine counterpart in the higher density region of the lensed distribution. In some cases, we also expect it to reduce the list of triggers requiring follow-up since some time delays have a zero probability under the lensed hypothesis. This effect is due to the fixed ordering of the images as well as our capacity to measure the difference in Morse factors unambiguously in most cases.
@@
To test the effectiveness of the method, we considered a simplified setup consisting of 50 lensed systems made of two or three super-threshold images and, at least, one sub-threshold image. For the super-threshold images, we ran joint parameter estimation to obtain their combined posteriors. To emulate an initial list of candidates, we generate sub-threshold events from a population model, construct posterior distributions following a Gaussian centred on the true values with widths scaled according to the events’ SNR, and compute the posterior overlap statistics between the chirp mass posterior for the unlensed counterpart and the super-threshold images. While not the same as the result from a genuine sub-threshold search system, this resembles other re-ranking methods proposed in the literature to follow-up on sub-threshold candidates. For every system considered, we compute our ranking statistics based on the measured time delay for the sub-threshold triggers. The final ranking is the product between the posterior overlap and our statistics. When doing this, we effectively reduce the number of triggers to follow-up, with an average factor improvement of 1.2, but the trigger list is more than halved when three super-threshold images are observed. Moreover, we have improved rankings in all cases. This demonstrates our method can improve the trigger list one would pass to more expensive follow-up searches.
@@
We note that, in principle, if more than one super-threshold image is identified, the joint posteriors for the two first images can be used to make the template bank and evaluate the sky overlap if desired. While this would probably not impact much the template bank itself, the sky overlap condition would probably become more stringent since joint parameter estimation effectively reduces the observed sky area (Janquart et al., 2021a). Moreover, sky overlap is complementary to our approach for reducing strong lensing detection candidates, and therefore using both the ranking based on sky map overlap and on the time delay compatibility would likely further reduce the list of candidates requiring follow-up. Finally, let us mention that if a pair of super-threshold events is detected, predicting and observing a third and/or fourth sub-threshold counterpart would strengthen the case for lensing.
@@
More testing and development are needed to use the method in realistic scenarios. For example, one could improve the selection of the measured relative lensing parameters to evaluate the $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ ratio, accounting for their full distribution and not their median values. In addition, one could verify the compatibility with other ranking and vetoing methods proposed in the literature, such as the ones used in Wong et al. (2021) and Goyal et al. (2023), for example. Ultimately, these methodologies should be tested on realistic sub-threshold search results, which is beyond the scope of this paper.
@@
In the end, this work shows that one can use the observed relative lensing parameters for a super-threshold lensed doublet or triplet to build a probability distribution for the time delay of a possible sub-threshold counterpart—predicting where such events should be situated in the data stream. Using a simplified setup, we showed that using this distribution to rank sub-threshold triggers is an effective approach because it generally promotes the genuine counterpart and reduces the list of triggers requiring more extensive follow-up. Finally, this method is also complementary to other ranking methods proposed in the literature, such as using the sky overlap between events. We hope that our work gives further motivation to the use of lensing time delay probability distributions in strong lensing searches as a mean to reduce the number of triggers requiring more expensive follow-up analyses.
@@ Acknowledgements
The data underlying this article will be shared on reasonable request to the corresponding authors.
@@ Appendix A Breakdown of the Probability Distributions for Different Sub-cases
In this appendix, we show a case-by-case breakdown of the probability distributions for the different sub-cases explained in Sec. 4 and Table 1. The total conditioned KDEs (same as the ones in Sec. 4) are shown on the left, and the corresponding breakdowns are shown on the right.
[Uncaptioned image][Uncaptioned image]
@@
[Uncaptioned image][Uncaptioned image]
@@
[Uncaptioned image][Uncaptioned image]
@@
[Uncaptioned image][Uncaptioned image]
@@
[Uncaptioned image][Uncaptioned image]
@@
[Uncaptioned image][Uncaptioned image]
@@
Appendix B Parametes of models used for simulation
The simulation was done by first sampling the lenses and BBHs from existing models. The sampling methods and the models are detailed in the documentation of ler. The parameters used to sample the BBHs population are listed below.
Parameter Value
$\lambda_{\text{peak}}$ $0.10$
$\alpha$ $2.63$
$\beta$ $1.26$
$\mu_{m}$ $33.07\ M_{\odot}$
$\sigma_{m}$ $5.69\ M_{\odot}$
$m_{\text{max}}$ $86.22\ M_{\odot}$
$m_{\text{min}}$ $4.59\ M_{\odot}$
$\delta_{m}$ $4.82\ M_{\odot}$
Table 1: BBH mass distributions model parameters, based on the GWTC-2 Population results (Abbott et al., 2021c).
Parameter Value
$\lambda_{z}$ $0.563$
$a$ $2.906$
$b$ $0.0158$
$c$ $0.58$
$\mu_{z}$ $1.1375$
$\sigma_{z}$ $0.8665$
Table 2: Parameters for $dV_{c}/dz$, the differential comoving volume (Wierda et al., 2021). This gives the redshift distribution of BBHs.